Status Report

NOAA SATOPS Morning Report: Thursday, June 07, 2012

By SpaceRef Editor
June 7, 2012
Filed under ,

SATOPS Morning Report: Thursday, June 07, 2012

Customers or users who want to call to ask about problems or report outages can use the following contacts:
– For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
– To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: [email protected]

DOY 159

GOES
GOES-12 DOY158: Six instrument cycle slips reported: 1033Z, 1207Z, 1242Z, 1324Z, 1325Z, and 1411Z.

GOES-12 Imager instrument electronics swap to side B executed 158/16:15Z. The 1627Z and 1630Z images were missed as a result. No cycle slips reported since the swap.

All other GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
All POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
All DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All other JASON-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
Gini NPORT 6 West products were not received due to equipment problems; 1445Z, 1500Z, 1530Z, 1545Z, 1600Z. Components were stopped and restarted to correct the problem.

Cape Ferguson data still not being received.

Cloud Data from Offutt lost from 158/2100Z until 159/0100Z.

Questions or comments to [email protected]

SpaceRef staff editor.